# Microcanonical simulated annealing: Massively parallel Monte Carlo simulations with sporadic random-number generation

### v2

@@ II.8 Putting all pieces together: the algorithms we tested
When applied, the walker-wandering part of the algorithm occurred once every timestep (the parity of the time index $t$ defines which of the two wandering operations is used, recall Sec. II.6). In the case with just $\gamma=1$ walker per spin, we employed forward-X wandering. In the case with $\gamma=6$ walkers per spin, we used
@@
In all cases, the refresh of the daemons/walkers took place at timesteps $t_{s}=2^{s}$, as specified in Sec. II.7. Observe that our definition of timestep ensures that the number of sweeps that a particular subfamily of daemons/walkers, $\alpha$, completes between consecutive refreshing steps is independent of $\gamma$.
II.8.1 The algorithm in the form of a procedure
@@
      If the $\{n_{\bm{x},\alpha}\}$ are walkers, carry out the part of the wandering cycle dictated by the parity of $t$, see Sect. II.6 (for daemons, do nothing).
@@
III Results
Figure 2: Comparison of Metropolis and MicSA dynamics with 6 walkers in the spin-glass phase at $T\!=\!0.7\approx 0.64T_{\mathrm{c}}$. a) Time evolution of the difference of the energy at time $t$ and the equilibrium value obtained through a fit to the canonical data. The solid yellow line is for the best fit described in Table 2. b) The coherence length, $\xi_{12}$, as a function of the simulation time for both models. The solid line is again for the best fit described in Table 2. c) Coherence length, $\xi_{12}$, as a function of the ratio between the coherence lengths $\xi_{23}$ and $\xi_{12}$. d) Difference between the results from the two algorithms for the energy, Eq. (22), as a function of the simulation time. e) As in d) but for the coherence length. To compute these differences, we have considered the best fit to our canonical data —solid lines in a) and b)— and the rescaling coefficient $k^{*}=6.034$ (see Appendix A for more information on the estimation of $k^{*}$). Error bars are smaller than data points. For the sake of clarity points are plotted at regular time intervals in logarithmic scale.
@@ III Results
$T$ $\gamma$ $k^{*}$ $\Delta k^{*}$ $[\xi_{\mathrm{min}},\xi_{\mathrm{max}}]$ $\chi^{2}/\mathrm{d.o.f.}$
@@
The form of the interpolating functions varies very much with temperature, as explained in Appendix A, where details on the above fits are provided (in particular, for $T\leq T_{\mathrm{c}}$ it is simpler to find $t$ as a function of $\xi_{12}$ than the other way around). The reader should take neither $F^{\xi}_{\text{can}}(T,t)$ nor $F^{e}_{\text{can}}(T,t)$ too far, because they are mere interpolating devices that are not good for extrapolations. Among other problems, these functions represent the results of a single sample, including times when —for $T\geq T_{\mathrm{c}}$— the condition $L\gg\xi(T,t)$ is no longer met.
@@
with the rescaling factor $k$ as the only fitting parameter. The best-fit rescaling parameter is named $k^{*}$ in Table 1 and in Figs. 23, and 4. In the figures, we have used for the energy the same value of $k^{*}$ computed with $\xi_{12}$. Each of the different variants of the MicSA algorithm will have its own $k^{*}$ that slightly depends on the temperature. For further details on the computation of $k^{*}$, see Fig. 5 and Appendix A.
Figure 5: Estimation of the best time shift $k^{*}$ at $T_{\mathrm{c}}$. In the main panel, the difference between the logarithm of the original measured time in the MicSA data and the interpolating function $f(k^{*}t)$ as a function of the coherence length $\xi_{12}$. The horizontal solid line at zero indicates the interval of the fit used to estimate $k^{*}$ (see Table 1 for details). Inset: $\log(t)$ as a function of $\xi_{12}$ for both algorithms. Solid lines indicate the best fit for our data. In the case of the MicSA data, the only fitting parameter is $k$, while the rest of the parameters $\Theta=\{a_{0},\,a_{1},\,z\}$ come from the fit to the canonical data described in Appendix A.
Table 2: Results of the different fits to our canonical data shown in Figs. 14. These fits correspond to the interpolation used to estimate $\delta_{e}(t,k)$ and $\delta_{\xi}(t,k)$ [see Eqs. (22) and (23)] explained in Sec. III.
$\mathcal{O}$ $T$ $F(t)$ $[t_{\mathrm{min}},t_{\mathrm{max}}]$ Fit $\chi^{2}/\mathrm{d.o.f.}$
[…]
Ideally speaking, we would hope $k^{*}$ to be close to $\gamma$, because in a single step of the microcanonical simulated annealing algorithm, every spin is updated $\gamma$ times (to be compared with just one spin update per canonical timestep). A direct inspection of Table 1 reveals that only the variants of the algorithm with walkers get close to this ideal. When one works with daemons rather than walkers, $k^{*}$ turns out to be significantly smaller than $\gamma$. Yet, as we shall discuss in the Conclusions, in order to choose the most convenient algorithmic variant, one should also take into account the computational cost of the different parts of the algorithm (see Appendix B).
@@ IV Conclusions
Our CUDA implementation demonstrates the parallel-friendly nature of the algorithm. Our code, see Appendix B, is publicly available and puts in the hands of a general user a computing performance that could be previously delivered only by the Janus II custom-built computer  [^3] In the context of quantum spin glasses, performance competitive with Janus II has been already obtained on GPUs [25].. Our program may run both standard Metropolis and MicSA, with the MicSA version taking half the time (in fact, in our Metropolis implementation, random-number generation takes about 75% of the simulation time). A standard metric corresponds to the number of picoseconds per spin update. Our MicSA program runs on an H100 GPU at 0.4 ps/spin-update, versus the 3.17 ps/spin-update on a single Janus II FPGA. The above two numbers should, however, be compared with care for two reasons. First, the computing power on the GPU is spread over 128 replicas, while on Janus II updates are concentrated on only two replicas. Hence, for a high-accuracy simulation spanning a limited time range (say up to $t\sim 10^{9}$), the GPU is faster. If, however, one needs to reach times $t\sim 10^{11}$ or larger (which is not uncommon [6]), Janus II still fares better (the GPU might make more spin updates per second by running many replicas, but would need a wall-clock time of many months to finish each simulation). Second, we are, somewhat unfairly, comparing 2025 high-end GPUs with the 2013 Virtex 7 FPGAs that Janus II is made of.
@@
The performance of MicSA for system sizes $L$ small enough to be brought to thermal equilibrium also in the spin-glass phase, $T<T_{\mathrm{c}}$, needs to be ascertained. In particular, MicSA needs to be combined with Parallel Tempering [65]. Although we do not foresee difficulties of principle, this algorithmic combination is yet to be tested.
@@
The performance of MicSA should also be investigated in problems where the elementary excitation (which is a spin flip in our case) will cause a continuously distributed energy change. The simplest example would be, perhaps, a spin glass with Gaussian-distributed couplings. Probably choosing the daemons/walkers as continuous, exponentially distributed random variables will fully solve the problem, but explicit tests are needed.
@@ Appendix A Computation of the best time shift
As we have explained in Sec. III, to estimate the best value of $k^{*}$, we use the coherence length $\xi_{12}$. A similar analysis can be performed with the energy, but, as it is clear from panel d) in Fig 2, the energy in the MicSA algorithm does not decrease to its equilibrium value monotonically. Instead, for short times, it is sensitive to the daemons’ refreshing process. This behavior makes it more challenging to estimate $k^{*}$ from these data. Nonetheless, the relation between the coherence length and time in the spin-glass phase, which can be extended to the critical point, offers a straightforward way to estimate $k^{*}$.
@@ Appendix B The CUDA implementation
In the following, we provide basic technical information about the CUDA implementation of the MicSA algorithm described in Sec. II applied to the three-dimensional Ising spin glass. The code implements a two-level parallelism: we use multispin coding to pack 64 spins in 64-bit words (long long int) and then CUDA threads for the concurrent update of even and odd spins. Multiple replicas and samples can be simulated in a single execution; the actual number of both depends on the resources (mainly memory) available on the GPU (we successfully tested up to 512 replicas). The code also implements the standard Metropolis algorithm using random numbers generated on the GPU with the procedure described in Ref. [25]. The refresh of the daemons/walkers is carried out on the CPU asynchronously to the microcanonical update (we recall that the refresh takes place according to an exponential schedule) using CPU threads. The walking procedure and the microcanonical update are implemented in two separate CUDA kernels to increase flexibility. The main reason is that the microcanonical update is carried out separately for each daemon/walker, whereas the walking procedure is completed for all the daemons in a single kernel execution.
@@
Acknowledgements.
We thank Martin Hasenbusch for calling our attention to Refs. [33, 34, 35, 36] and for kindly explaining to us some of the details in these works. This work was partially supported by MICIU/AEI/10.13039/501100011033 (Spain) and by “ERDF: A way of making Europe” (European Union) through grants no. PID2022-136374NB-C21, PID2020-112936GB-I00 and RED2022-134244-T; by Junta de Extremadura (Spain) through grant no. GR24022; and by funding from the 2021 first FIS (Fondo Italiano per la Scienza) funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) from Italian MUR (Ministry of University and Research). We acknowledge the use of the CESAR computational resources at the BIFI Institute (University of Zaragoza). IG, MB and FRT acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP E83C22003230001, CUP B93C22000620006 and CUP B83C22002940006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.
Data availability
Our CUDA code and the data required to reproduce our figures are publicly available at [66].

### v3

@@ 2.8 Putting all pieces together: the algorithms we tested
When applied, the walker-wandering part of the algorithm occurred once every timestep (the parity of the time index $t$ defines which of the two wandering operations is used, recall Sec. 2.6). In the case with just $\gamma=1$ walker per spin, we employed forward-X wandering. In the case with $\gamma=6$ walkers per spin, we used
@@
In all cases, the refresh of the daemons/walkers took place at timesteps $t_{s}=2^{s}$, as specified in Sec. 2.7. Observe that our definition of timestep ensures that the number of sweeps that a particular subfamily of daemons/walkers, $\alpha$, completes between consecutive refreshing steps is independent of $\gamma$.
2.8.1 The algorithm in the form of a procedure
@@
      If the $\{n_{\bm{x},\alpha}\}$ are walkers, carry out the part of the wandering cycle dictated by the parity of $t$, see Sect. 2.6 (for daemons, do nothing).
@@
3 Results
Figure 2: Comparison of Metropolis and MicSA dynamics with 6 walkers in the spin-glass phase at $T\!=\!0.7\approx 0.64T_{\mathrm{c}}$. a) Time evolution of the difference of the energy at time $t$ and the equilibrium value obtained through a fit to the canonical data. The solid yellow line is for the best fit described in Table 2. b) The coherence length, $\xi_{12}$, as a function of the simulation time for both models. The solid line is again for the best fit described in Table 2. c) Coherence length, $\xi_{12}$, as a function of the ratio between the coherence lengths $\xi_{23}$ and $\xi_{12}$. d) Difference between the results from the two algorithms for the energy, Eq. (21), as a function of the simulation time. e) As in d) but for the coherence length. To compute these differences, we have considered the best fit to our canonical data —solid lines in a) and b)— and the rescaling coefficient $k^{*}=6.034$ (see A for more information on the estimation of $k^{*}$). Error bars are smaller than data points. For the sake of clarity points are plotted at regular time intervals in logarithmic scale.
@@ 3 Results
$T$ $\gamma$ $k^{*}$ $\Delta k^{*}$ $\xi_{\mathrm{min}},\xi_{\mathrm{max}}]$ $\chi^{2}/\mathrm{d.o.f.}$
@@
The form of the interpolating functions varies very much with temperature, as explained in A, where details on the above fits are provided (in particular, for $T\leq T_{\mathrm{c}}$ it is simpler to find $t$ as a function of $\xi_{12}$ than the other way around). The reader should take neither $F^{\xi}_{\text{can}}(T,t)$ nor $F^{e}_{\text{can}}(T,t)$ too far, because they are mere interpolating devices that are not good for extrapolations. Among other problems, these functions represent the results of a single sample, including times when —for $T\geq T_{\mathrm{c}}$— the condition $L\gg\xi(T,t)$ is no longer met.
@@
with the rescaling factor $k$ as the only fitting parameter. The best-fit rescaling parameter is named $k^{*}$ in Table 1 and in Figs. 23, and 4. In the figures, we have used for the energy the same value of $k^{*}$ computed with $\xi_{12}$. Each of the different variants of the MicSA algorithm will have its own $k^{*}$ that slightly depends on the temperature. For further details on the computation of $k^{*}$, see Fig. 5 and A.
Figure 5: Estimation of the best time shift $k^{*}$ at $T_{\mathrm{c}}$. In the main panel, the difference between the logarithm of the original measured time in the MicSA data and the interpolating function $f(k^{*}t)$ as a function of the coherence length $\xi_{12}$. The horizontal solid line at zero indicates the interval of the fit used to estimate $k^{*}$ (see Table 1 for details). Inset: $\log(t)$ as a function of $\xi_{12}$ for both algorithms. Solid lines indicate the best fit for our data. In the case of the MicSA data, the only fitting parameter is $k$, while the rest of the parameters $\Theta=\{a_{0},\,a_{1},\,z\}$ come from the fit to the canonical data described in A.
Table 2: Results of the different fits to our canonical data shown in Figs. 14. These fits correspond to the interpolation used to estimate $\delta_{e}(t,k)$ and $\delta_{\xi}(t,k)$ [see Eqs. (21) and (22)] explained in Sec. 3.
$\mathcal{O}$ $T$ $F(t)$ $t_{\mathrm{min}},t_{\mathrm{max}}]$ Fit $\chi^{2}/\mathrm{d.o.f.}$
$e$ $0.7$ $e_{\infty}+a_{1}t^{-z_{1}}+a_{2}t^{-z_{2}}$ $1.7\times 10^{2},*]$ $\begin{aligned} e_{\infty}=-1.7708874(19),\,a_{1}=0.0884(7),a_{2}=0.1060(3),\\[-4.0pt] z_{1}=0.2235(5),\,z_{2}=0.445(3)\hskip 17.00024pt\hskip 17.00024pt\qquad\end{aligned}$ $242.4/248$
$e$ $T_{\mathrm{c}}$ $e_{\infty}+a_{1}t^{-z_{1}}$ $10^{4},*]$ $e_{\infty}=-1.7022747(19),\,a_{1}=0.1325(12),z_{1}=0.3820(9)$ $152.4/149$
$e$ $1.4$ $e_{\infty}+a_{1}t^{-z_{1}}$ $10^{4},*]$ $e_{\infty}=-1.607026(4),\,a_{1}=2.5(9),z_{1}=0.90(4)$ $91.7/83$
$\xi_{12}$ $0.7$ $a_{0}+a_{1}t^{z}_{1}+a_{2}\log(t)$ $1.4\times 10^{2},*]$ $a_{0}=-2.71(3),\,a_{1}=4.10(3),a_{2}=-0.1179(14),z=0.05930(17)$ $197.6/235$
$\xi_{12}$ $T_{\mathrm{c}}$ $a_{0}+a_{1}t^{z}_{1}$ $10^{3},10^{9}]$ $a_{0}=0.159(8),\,a_{1}=1.158(4),z=0.1543(2)$ $126.4/157$
$\xi_{12}$ $1.4$ $C\left(1-\sum_{i=0}^{7}a_{i}\mathrm{e}^{-tb^{i}/\tau}\right)$ $10^{3},*]$ $\begin{aligned} C=9.29(2),\,\tau=8(3)\times 10^{6},b=4.0(4),\hskip 17.00024pt\hskip 17.00024pt\\[-4.0pt] a_{0}=0.007(4),\,a_{1}=0,\,a_{2}=0.11(3),a_{3}=0.166(17),\qquad\\[-4.0pt] a_{4}=0.136(9),\,a_{5}=0.123(9),\,a_{6}=0.095(7),a_{7}=0.069(8)\end{aligned}$ $49.4/104$
Ideally speaking, we would hope $k^{*}$ to be close to $\gamma$, because in a single step of the microcanonical simulated annealing algorithm, every spin is updated $\gamma$ times (to be compared with just one spin update per canonical timestep). A direct inspection of Table 1 reveals that only the variants of the algorithm with walkers get close to this ideal. When one works with daemons rather than walkers, $k^{*}$ turns out to be significantly smaller than $\gamma$. Yet, as we shall discuss in the Conclusions, in order to choose the most convenient algorithmic variant, one should also take into account the computational cost of the different parts of the algorithm (see B).
@@ 4 Conclusions
Our CUDA implementation demonstrates the parallel-friendly nature of the algorithm. Our code, see B, is publicly available and puts in the hands of a general user a computing performance that could be previously delivered only by the Janus II custom-built computer. [^2] In the context of quantum spin glasses, performance competitive with Janus II has been already obtained on GPUs [25]. Our program may run both standard Metropolis and MicSA, with the MicSA version taking half the time (in fact, in our Metropolis implementation, random-number generation takes about 75% of the simulation time). A standard metric corresponds to the number of picoseconds per spin update. Our MicSA program runs on an H100 GPU at 0.4 ps/spin-update, versus the 3.17 ps/spin-update on a single Janus II FPGA. The above two numbers should, however, be compared with care for two reasons. First, the computing power on the GPU is spread over 128 replicas, while on Janus II updates are concentrated on only two replicas. Hence, for a high-accuracy simulation spanning a limited time range (say up to $t\sim 10^{9}$), the GPU is faster. If, however, one needs to reach times $t\sim 10^{11}$ or larger (which is not uncommon [6]), Janus II still fares better (the GPU might make more spin updates per second by running many replicas, but would need a wall-clock time of many months to finish each simulation). Second, we are, somewhat unfairly, comparing 2025 high-end GPUs with the 2013 Virtex 7 FPGAs that Janus II is made of.
@@
The performance of MicSA for system sizes $L$ small enough to be brought to thermal equilibrium also in the spin-glass phase, $T<T_{\mathrm{c}}$, needs to be ascertained. In particular, MicSA needs to be combined with Parallel Tempering [63]. We are already testing this algorithm combination for small systems, see C and Ref. [64].
@@
The performance of MicSA should be investigated in problems where the elementary excitation (which is a spin flip in our case) will cause a continuously distributed energy change. The simplest example would be, perhaps, a spin glass with Gaussian-distributed couplings. As C shows, choosing the daemons/walkers as continuous, exponentially distributed random variables fully solves the problem.
@@
Data availability
Our CUDA code and the data required to reproduce the figures are publicly available at [65].
Acknowledgments
We thank Martin Hasenbusch for calling our attention to Refs. [33, 34, 35, 36] and for kindly explaining to us some of the details in these works.
This work was partially supported by Ministerio de Ciencia, Innovación y Universidades (Spain) and by the European Regional Development Fund (MCIU/AEI/10.13039/501100011033/FEDER, UE) through grants no. PID2022-136374NB-C21, PID2024-156352NB-I00, PID2024-158623NB-C22 and RED2022-134244-T; by Junta de Extremadura (Spain) through grant no. GR24022; and by funding from the 2021 first FIS (Fondo Italiano per la Scienza) funding scheme (FIS783 - SMaC - Statistical Mechanics and Complexity) from Italian MUR (Ministry of University and Research). We acknowledge the use of the CESAR computational resources at the BIFI Institute (University of Zaragoza). IG, MB and FRT acknowledge the support of the National Center for HPC, Big Data and Quantum Computing, Project CN_00000013 – CUP E83C22003230001, CUP B93C22000620006 and CUP B83C22002940006, Mission 4 Component 2 Investment 1.4, funded by the European Union – NextGenerationEU.
@@ Appendix A Computation of the best time shift
As we have explained in Sec. 3, to estimate the best value of $k^{*}$, we use the coherence length $\xi_{12}$. A similar analysis can be performed with the energy, but, as it is clear from panel d) in Fig 2, the energy in the MicSA algorithm does not decrease to its equilibrium value monotonically. Instead, for short times, it is sensitive to the daemons’ refreshing process. This behavior makes it more challenging to estimate $k^{*}$ from these data. Nonetheless, the relation between the coherence length and time in the spin-glass phase, which can be extended to the critical point, offers a straightforward way to estimate $k^{*}$.
@@ Appendix B The CUDA implementation
In the following, we provide basic technical information about the CUDA implementation of the MicSA algorithm described in Sec. 2 applied to the three-dimensional Ising spin glass. The code implements a two-level parallelism: we use multispin coding to pack 64 spins in 64-bit words (long long int) and then CUDA threads for the concurrent update of even and odd spins. Multiple replicas and samples can be simulated in a single execution; the actual number of both depends on the resources (mainly memory) available on the GPU (we successfully tested up to 512 replicas). The code also implements the standard Metropolis algorithm using random numbers generated on the GPU with the procedure described in Ref. [25]. The refresh of the daemons/walkers is carried out on the CPU asynchronously to the microcanonical update (we recall that the refresh takes place according to an exponential schedule) using CPU threads. The walking procedure and the microcanonical update are implemented in two separate CUDA kernels to increase flexibility. The main reason is that the microcanonical update is carried out separately for each daemon/walker, whereas the walking procedure is completed for all the daemons in a single kernel execution.
@@
Appendix C A Parallel Tempering example (walkers in equilibrium simulations)
Figure 6: Overlap distribution of some selected temperatures. The curves for each method are obtained by simulating 200 times the same sample for 250000 EMCS (half of which have been used for prethermalization, meaning that we keep data only from the second half of the simulation) and later averaging over the 200 curves obtained. The simulation parameters are $L=10$, $N_{R}=2$, $N_{\beta}=40$, temperatures uniformly distributed between $T_{\min}=1/\beta_{\max}=0.2$ and $T_{\max}=1/\beta_{\min}=2$.
In this section, we provide an example of the possible use of the walkers for a simulation at equilibrium, which radically differs from the off-equilibrium dynamics studied in the main text. Bringing a spin-glass sample to equilibrium is possible only if the linear size of the system is tiny (particularly at low temperatures). Furthermore, special simulation methods, such as Parallel Tempering [63], are mandatory. The particular problem that we shall investigate here is a simple variant of Eq. (1) in which the couplings are drawn from a Gaussian distribution $J\sim\mathcal{N}(0,1)$. The rationale for choosing Gaussian-distributed couplings is that we wish to test our methods in a demanding situation, namely at the very low temperature $T=0.2\approx 0.21T_{\mathrm{c}}$ ($T_{\mathrm{c}}$ is the critical temperature). Indeed, continuously distributed couplings are usually preferable for very-low-temperature studies.
Specifically, we shall study in detail a single sample of a 3D system with lattice linear size $L=10$. For this sample, the walkers method reproduces the equilibrium values of the energy per spin [i.e., the spin Hamiltonian in Eq. (1) divided by the number of spins $N$] and of the overlap between the instantaneous value of two replicas
$q=\frac{1}{N}\sum_{\bm{x}}s_{\bm{x}}^{(a)}s_{\bm{x}}^{(b)}.$ (26)
In order to do this, we adapted the walker method to the Parallel Tempering procedure enhanced with Houdayer moves [66], which has been proven to be (mildly) effective also for the equilibrium simulation of three-dimensional spin glasses [67], even though our schedule is a bit different [64]. We are then running $N_{R}=2$ replicas of $N_{\beta}=40$ clones with temperatures uniformly distributed between $T_{\min}=1/\beta_{\max}=0.2$ and $T_{\max}=1/\beta_{\min}=2$.
In this case, our method requires the use of continuous walker variables, distributed as in Eq. (5). Additionally, we do not need to limit the capacity of the walkers as in Sec. 2.5, but we can take a step back and use the elementary move described in Sec. 2.3.
Table 4: Measures of the equilibrium expectation values for $|q|$, $e_{\infty}$ as extracted from the simulations described in C, at the lowest temperature $T_{\text{min}}=0.2$. For each of the two algorithms, we carried out 200 independent runs. Errors are straightforwardly computed by considering the fluctuations between runs.
$|q|$ error $e_{\infty}$ error
Walkers 0.743 8 0.012 9 1.696 330 17-1.696\,330\,17 0.000 001 02
Metropolis 0.741 6 0.010 4 1.696 331 13-1.696\,331\,13 0.000 000 89
Difference 0.002 19 0.001 67 $\phantom{-}0.000\,009\,65$ 0.000 001 36
For each replica/clone, we simulate a system with one walker per spin ($\gamma=1$) that gets refreshed at each Elementary Monte Carlo Step (EMCS) as described below.
The Elementary Monte Carlo Step (EMCS) is then structured as follows:
  1. 1.
    For each clone and for each replica, perform a fixed-temperature move:
    1. (a)
      Generate then walkers according to Eq. (5).
    2. (b)
      Alternate a full-lattice sweep with a wandering of the walkers towards the foward-X direction. Repeat 10 times.
  2. 2.
    Perform Houdayer’s move.
  3. 3.
    Perform a Parallel Tempering step between the $N_{\beta}$ clones belonging to the same replica group.
Fig. 6 qualitatively shows how the probability distribution of $q$ computed in this manner reproduces the results of using a canonical Metropolis update instead of point 1 of the above algorithm.
As a quantitative test, we have also evaluated the average energy per spin and the average overlap (in absolute value) yielded by both simulations. The obtained values are summarized in Table 4, which displays good agreement between the two algorithms. Furthermore, from the ratio between the errors we can estimate the ratio of the autocorrelation integrated time for both algorithms [53]. We obtain for $|q|$ the ratio $\tau_{\text{walkers}}/\tau_{\text{Metropolis}}\simeq 1.7$, meaning that, in this setting, one EMCS with the walkers is roughly equivalent to $0.59$ steps with the usual Metropolis. We can understand which of the two ways is more effective only by measuring also the CPU time for one EMCS, which depends not only on hardware, but also on parameters of the simulation such as the linear size $L$ or the number of temperatures $N_{\beta}$ of the Parallel Tempering. [^3] For instance, our partially multi-spin coded program runs about twice as fast with walkers as with Metropolis when $L=18$ on an AMD EPYC 7763 64-Core Processor; the computational advantage of the walker algorithm seems to grow with $L$.